i am in desperate need of some help...i do not understand how to do this it confuses me...
it says "using f(x)=2x+1 find..
f(-1)=
f(3.2)
im not looking for the answers im lookin to learn how to solve these so that i can do it on my own...any help would be greatly appericiated..
thanks

Answer by venugopalramana(3286)

im not looking for the answers im lookin to learn how to solve these so that i can do it on my own...
GREAT THAT IS THE SPIRIT..YOU WILL DEFINITELY SUCCEED..LET ME EXPLAIN
using f(x)=2x+1 find..
f(-1)=
f(3.2)=
f(x) means a function of x.that is you can give any permissible values to x (permissible so that the given function has some meaning..you will know later about domain as it is called)and substitute that value of x in r.h.s. of the equation given for f(x) to get the value of the function at that value of x ..so f(a) means ,put a in place of x wherever it occurs on the rhs and simplify
here f(x)=2x+1..we have to find f(-1)..that is put x=-1 on the r.h.s...so we get f(-1)=2(-1)+1=-2+1=-1..now you can be true to your word and continue with other example!!!!